# JANET: Joint Adaptive predictioN-region Estimation for Time-series

### v1

@@ Appendix B Training Details
B.1 Single time series experiments
@@
B.2 Particle experiments
We take the same experimental setup as Sun and Yu (2024) and use a 1-layer sequence-to-sequence LSTM network (EncDec) where the encoder has a single LSTM layer with embedding size 24 and the decoder has a single LSTM layer with embedding size 24 and a linear layer. We also fit a 1-layer RNN with a single LSTM layer (RNN) with embedding size 24, followed by a linear layer. We train the model for 150 epochs and set batch size to 150.
@@ B.2 Particle experiments
B.3 UK COVID experiments
We take the same architectures from the particle experiments and apply them to the same COVID data as Stankeviciute et al. (2021). We train these models for 200 epochs with embedding sizes of 128 and set batch size of 64.
@@ B.3 UK COVID experiments
B.4 Complete Monte Carlo results
We visualise the results of these experiments in Figure 3. For complete numerical results see Tables 3-8. We find that our JANET methods are comparable or better with respect to coverage and interval widths against the baselines.

### v2

@@ Appendix D Training Details
D.1 Single time series experiments
@@
D.2 Particle experiments
We take the same experimental setup as [38] and use a 1-layer sequence-to-sequence LSTM network (EncDec) where the encoder has a single LSTM layer with embedding size 24 and the decoder has a single LSTM layer with embedding size 24 and a linear layer. We also fit a 1-layer RNN with a single LSTM layer (RNN) with embedding size 24, followed by a linear layer. We train the model for 150 epochs and set batch size to 150.
@@ D.2 Particle experiments
D.3 UK COVID experiments
We take the same architectures from the particle experiments and apply them to the same COVID data as [35]. We train these models for 200 epochs with embedding sizes of 128 and set batch size of 64.
@@ D.3 UK COVID experiments
D.4 Complete Monte Carlo results
We visualise the results of these experiments in Figure 4. For complete numerical results see Tables 8-13. We find that our JANET methods are comparable or better with respect to coverage and interval widths against the baselines.